# Informativity Conditions for Multiple Signals: Properties, Experimental Design, and Applications (extended version)

### v3

@@ A. Proof of Theorem 1
$~~~-~\text{dim}[\text{im}\,{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})^{T}}\cap\text{im}({{\bf{1}}_{p}}\otimes{I_{{T_{0}}-L+1}})^{\perp}].$
@@
if $\text{im}\,{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})^{T}}\cap\text{leftker}\,{\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$, i.e., $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE.
@@
Similarly to the above, if additionally, $\text{im}\,{{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{\bar{p}}})}^{T}}\cap\text{leftker}\,{\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$, then $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE.
3). i). If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE, then $T_{0}=T_{1}=...=T_{\bar{p}}$ and ${H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}{H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{T}>0$, i.e., ${M_{1}}+\sum\limits_{i=1}^{\bar{p}}{\left({{\alpha_{i}}{H_{L}}({z_{i}^{[0,{T_{0}}-1]}})\sum\limits_{i\neq j=1}^{\bar{p}}{{\alpha_{j}}H_{L}^{T}({z_{j}^{[0,{T_{0}}-1]}})}}\right)}>0$. Similar to the proof of conclusion 1), we can deduce $M_{1}>0$, and thus $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also MCPE.
@@
B. Proof of Lemma 1
Proof: We begin by demonstrating that
@@ B. Proof of Lemma 1
C. Proof of Lemma 3
According to condition ii), there must exists some invertible matrix sequence $M(k)$ such that
$2I-{z^{T}}(k)z(k)=M(k){M^{T}}(k)$
for all $k\geq 0$. Let $A(k)=I-z(k)z^{T}(k)$ and $N(k)=z(k)M(k)$, we have
$A^{T}(k)A(k)-I=-N(k){N^{T}}(k)\leq 0.$
Consider the Lyapunov function $V(k)=x^{T}(k)x(k)$, we can compute the time difference of $V(k)$ as
$V(k+1)-V(k)$ $=-x^{T}(k)(A^{T}(k)A(k)-I)x(k)$
$\leq 0,$
which implies that the equilibrium $x_{e}=0$ is stable in the sense of Lyapunov. Recall the result (Ioannou and Fidan, 2006, Th. A.12.22) of exponential stability, if, further, the pair $(A(k)-K(k)N^{T}(k),N(k))$ is uniformly completely observable (UCO) for any bounded $K(k)$, then $x_{e}=0$ is e.s. Let us choose
$K(k)=-z(k){M^{T}}(k)^{-1},$
which is bounded due to condition i). Then we have $A(k)-K(k)N^{T}(k)=I$. Therefore, it is only necessary to prove the fact that the pair $(I,N(k))$ is UCO.
If $z_{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$, we have $H_{L}(z_{[k,k+l-1]})H^{T}_{L}(z_{[k,k+l-1]})>0$. Hence, there exists a constant $\beta_{0}>0$ such that
$H_{1}(z_{[k,k+l-1]})H^{T}_{1}(z_{[k,k+l-1]})$ $=\sum\limits_{t=0}^{l-1}{z(k+t)z^{T}(k+t)}$
$\geq{\beta_{0}}I.$ (29)
By (C. Proof of Lemma 3) and condition i), we know that
${\beta_{0}}I\leq\sum\limits_{t=0}^{l-1}{z(k+t){z^{T}}(k+t)}\leq{\beta_{1}}I$ (30)
for some $\beta_{1}\geq\beta_{0}>0$. In addtion, since $N(k){N^{T}}(k)=z(k)(2I-{z^{T}}(k)z(k)){z^{T}}(k)$, one has
$(2-z_{m}^{2})\sum\limits_{t=0}^{l-1}{z(k+t){z^{T}}(k+t)}\leq\sum\limits_{t=0}^{l-1}{N(k+t){N^{T}}(k+t)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\leq 2\sum\limits_{t=0}^{l-1}{z(k+t){z^{T}}(k+t)}.$ (31)
Combining (30) and (C. Proof of Lemma 3), we have
${\beta_{0}}(2-z_{m}^{2})I$ $\leq\sum\limits_{t=0}^{l-1}{N(k+t){N^{T}}(k+t)}\leq 2{\beta_{1}}I,$ (32)
which implies that $(I,N(k))$ is UCO and the equilibrium state $x_{e}=0$ is e.s. $\blacksquare$

### v4

@@ A. Proof of Theorem 1
$~~~-~\text{dim}[\text{im}\,{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})^{\top}}\cap\text{im}({{\bf{1}}_{p}}\otimes{I_{{T_{0}}-L+1}})^{\perp}].$
@@
if $\text{im}\,{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})^{\top}}\cap\text{leftker}\,{\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$, i.e., $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE.
@@
Similarly to the above, if additionally, $\text{im}\,{{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{\bar{p}}})}^{\top}}\cap\text{leftker}\,{\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$, then $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE.
3). i). If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE, then $T_{0}=T_{1}=...=T_{\bar{p}}$ and ${H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}{H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{\top}>0$, i.e., ${M_{1}}+\sum\limits_{i=1}^{\bar{p}}{\left({{\alpha_{i}}{H_{L}}({z_{i}^{[0,{T_{0}}-1]}})\sum\limits_{i\neq j=1}^{\bar{p}}{{\alpha_{j}}H_{L}^{\top}({z_{j}^{[0,{T_{0}}-1]}})}}\right)}>0$. Similar to the proof of conclusion 1), we can deduce $M_{1}>0$, and thus $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also MCPE.
@@
B. Proof of Lemma 3.1
We begin by demonstrating that